. .NEED ANSWER ASAP / ANSWER NEVER USED BEFORE, COMPLETELY NEW ANSWER PLEASE a.) Find the...

.

.NEED ANSWER ASAP / ANSWER NEVER USED BEFORE, COMPLETELY NEW ANSWER PLEASE

a.)

Find the following values, using the equations, and then work the problems using a financial calculator to check your answers. Disregard rounding differences. (Hint: If you are using a financial calculator, you can enter the known values and then press the appropriate key to find the unknown variable. Then, without clearing the TVM register, you can "override" the variable that changes by simply entering a new value for it and then pressing the key for the unknown variable to obtain the second answer. This procedure can be used in parts b and d, and in many other situations, to see how changes in input variables affect the output variable.) Do not round intermediate calculations. Round your answers to the nearest cent.

  1. An initial $200 compounded for 1 year at 7.5%.

    $  

  2. An initial $200 compounded for 2 years at 7.5%.

    $  

  3. The present value of $200 due in 1 year at a discount rate of 7.5%.

    $  

  4. The present value of $200 due in 2 years at a discount rate of 7.5%.

    $  

b.)

Use both the TVM equations and a financial calculator to find the following values. (Hint: If you are using a financial calculator, you can enter the known values and then press the appropriate key to find the unknown variable. Then, without clearing the TVM register, you can "override" the variable that changes by simply entering a new value for it and then pressing the key for the unknown variable to obtain the second answer. This procedure can be used in parts b and d, and in many other situations, to see how changes in input variables affect the output variable.) Do not round intermediate calculations. Round your answers to the nearest cent.

  1. An initial $600 compounded for 10 years at 4%.

    $  

  2. An initial $600 compounded for 10 years at 8%.

    $  

  3. The present value of $600 due in 10 years at a 4% discount rate.

    $  

  4. The present value of $600 due in 10 years at an 8% discount rate.

    $  
Answer #1

Solution:a)

FV = PV x (1 + r)^n

a) FV = 200 x (.01+ 075)^ 1 = 200 x (1.075)^1 = $215

b) FV = 200 x (.01+ 075)^2 = $231.12

C & d) Calculation of PV:

Cashflow (a) Discounting factor @ 7.5 years (b) PV (a x b)
200 For 1 year = 0.9302 $186
200 For 2 years = 1.7956 $ 359.12 r/o $360

Solution:b)

FV = PV x (1 + r)^n

a) FV = 600 x (.01+ 04)^ 10 = 600 x (1.04)^10 = $888

b) FV = 600 x (.01+ 08)^10 = 600 x (1.08)^10 = 1295.35

C & d) Calculation of PV:

Cash flow (a) Discounting factor for 10 years(b) PV (a x b)
600 @ 4% = 8.1109 $4867
600 @ 8% = 6.7109 $ 4026